### v4

# In this paper, we consider the problem whether there exist transcendental meromorphic functions which satisfy the functional equation

Project supported in part by the NSF of Shandong Province, China (No. ZR2019MA029),
the FRFCU(No.3016000842464005 ), the NSF of Shandong Province, China (No.ZR2014AM011) and the NSFC (No.11171184).
1. Introduction and main results
In this paper, we consider the problem whether
there exist transcendental meromorphic functions which satisfy the
functional equation
where is an integer. By meromorphic functions, we will always mean meromorphic functions in the complex plane
In 1985, W.K.Hayman proved the following result:
Theorem 1.
([19]) There do not exist non-constant meromorphic functions satisfying
(1) for
Later on, Gundersen [10] and Gundersen [11] gave examples of transcendental meromorphic functions satisfying (1) for the cases and For the cases it is known that there exist transcendental entire functions satisfying (1); see [10].
Theorem 1 can also be proved by using earlier results of Fujimoto [6, Corollary 6.4] or Green
[8, Lemma on p. 71]. For rational functions, Hayman [19] proved the next result:
Theorem 2.
([19]) There do not exist non-constant rational functions satisfying
(1) for
In 1971, Toda [34] proved the following result:
Theorem 3.
([34]) There do not exist non-constant entire functions satisfying
(1) for
Gundersen-Tohge [15] gave examples of transcendental entire functions satisfying (1) for It is known that there exist such examples for the cases see [16]. For polynomials, Newman-Slater [30] proved the next result:
Theorem 4.
([30]) There do not exist non-constant polynomials satisfying
(1) for
We mention that Hayman [19] showed that Theorems 1-4 can all be proved by using Cartanβs version of Nevanlinna Theory.
From the above discussion, one may ask the following questions:
Question 1.
([14] and [16, Question 3.1]) β whether or not there exist non-constant meromorphic functions satisfying for the cases and ?
Question 2.
([14] and [16, Question 3.3]) β whether or not there exist non-constant entire functions satisfying ?
Throughout the paper, we adopt the standard
notations etc., in the Nevanlinna theory of a meromorphic function We assume the readers are familiar with the fundamental results and the typical kind of arguments in Nevanlinna theory; see for example, in [2, 7, 18, 22, 29, 35, 36]. It will be convenient to let denote any set of positive real numbers of finite linear
measure, not necessarily the same set at each occurrence. Thus for a
non-constant meromorphic function we denote by any quantity
satisfying as where is the Nevanlinna characteristic of Following [17, p.466], we define the notion of the small function of a non-constant meromorphic function in the complex plane as follows: let be a non-constant meromorphic function. and let be also a meromorphic function in the complex plane such that for all possibly outside a set of finite linear measure. Then, the meromorphic function is called small compared to or slowly moving with respect to or a small function of for short.
Regarding Questions 1 and 2, we recall the following results from Ishizaki [21]:
Theorem 5.
([21, Proposition 5.1]).
Suppose that there exist transcendental meromorphic functions
satisfying the functional equation (1) for and suppose
that and are linearly independent. Then there exists a small function
with respect to and such that
where and in what follows, is the
Wronskian determinant of and
Theorem 6.
([21, Proposition 6.1]).
Suppose that there exist transcendental functions
satisfying the functional equation (1) for and suppose
that and are linearly independent. Then there exists a small function
with respect to and such that
For the existence of non-constant meromorphic solutions of a much more general three term functional equation, we recall the following result from [20] and Yang-Zhang [37, 38]:
Theorem 7.
([20, 37, 38])
If
then there do not exist any non-constant meromorphic functions
satisfying the functional equation where and are positive integers such that
We will prove the following result:
Theorem 8.
There do not exist transcendental meromorphic solutions satisfying the functional equation
In the same manner as in the proof of Theorem 8 in Section 3, we can get the following result:
Theorem 9.
There do not exist transcendental entire solutions satisfying the functional equation
By combining Theorems 2 and 4 with Theorems 8 and 9, we obtain the following results:
Corollary 1.
There do not exist non-constant meromorphic solutions satisfying the functional equation
Corollary 2.
There do not exist non-constant entire solutions satisfying the functional equation
Corollary 1 resolves Question 1 for and Corollary 2 resolves Question
2 completely.
2. Preliminaries
In this section, we will introduce some important lemmas to prove the main results in this paper. We first give the following notations: Let be a non-constant meromorphic function in the complex plane, and let be a point. Next we set
and
Similarly, we define
and
Let and be two non-constant meromorphic functions, and let
We recall the following results from Yi[39]:
Lemma 1.
([39]) The following items hold:
(i) If and then
(ii) If and then
(iii) If then
(iv) If and then
(v) If and then
(vi) If and then
Next we let be non-constant meromorphic functions satisfying the following functional equation
where and in what follows, and are positive integers such that Next we also set
The following Lemma 2-Lemma 12 were proved originally in the Chinese academic paper Yi-Yang[40]. For convenience of the readers, we provide their detailed proofs as follows:
Lemma 2.
Suppose that and are non-constant meromorphic functions satisfying (4). If and then
where and in what follows, and are defined as in (5), is any quantity such that as and Here is a subset with finite linear measure, and
Proof.
Suppose that Then, by (4) we have
which implies that
Noting that is a non-constant meromorphic function, by rewriting (4) we have
where By (6), (7), (8) and the second fundamental theorem we have
this together with and (5) implies a contradiction. Therefore, Similarly we have
and
β
Lemma 3.
Suppose that and are non-constant meromorphic functions satisfying (4). If and
where and are defined as in (5), then
and are linearly independent.
Proof.
Without loss of generality, we suppose that and Then, by the assumption of Lemma 3 we have
and Suppose that and are linearly dependent. Then, there exist three constants and satisfying say such that Combining this with (4) we have
Noting that and are non-constant meromorphic functions, we have by (9) that and Therefore, by (9)
and the second fundamental theorem we have
and
which together with and gives a contradiction.
β
The following result is due to Nevanlinna [28]:
Lemma 4.
([28])
Suppose that are linearly independent meromorphic functions
such that Then, for
where is the Wronskian determinant of and is defined in a similar way to Lemma 2.
Let be three non-constant meromorphic functions satisfying (4). Then, the Wronskian determinant of and is such that
By Lemmas 3 and 4 we can get the following result:
Lemma 5.
Let and If are three non-constant meromorphic functions satisfying (4), then
where is in (10) and is defined as in Lemma 2.
For proving the following lemma, we first introduce the following notations: Let be three non-constant meromorphic functions satisfying (4), and let be a point such that is either a zero or a pole of any one of the three meromorphic functions and Then, the Laurent developments of and at the point are
and
respectively, where and are nonzero constants, and are three integers such that Next we also let
where is in (10).
Let and be non-constant meromorphic functions, and let be a point. Next we define
By using Lemma 1, we will prove the next several results:
Lemma 6.
Suppose that and Then
Proof.
Since are non-constant meromorphic functions satisfying (4), we deduce by (11)- (13) and the assumption of Lemma 6 that at least one of is equal to hence
We consider the following two cases:
Case 1. β Suppose that there exists only one of of is a positive integer, say Then, by (10) and Lemma 1(i) we have
By (14) and (16) we have
which implies that the conclusion of Lemma 6 is valid.
Case 2. β Suppose that exactly two of are positive integers, say and Then, by (10) and Lemma 1(ii) we have
By (14) and (17) we have
which reveals the conclusion of Lemma 6.
β
Lemma 7.
Suppose that and Then
Proof.
Since are non-constant meromorphic functions satisfying (4), we deduce by (11)-(13) and the assumption of Lemma 7 that only one of is a positive integer, the other two of are negative integers, say and This together with (10) and Lemma 1(vi) gives
By (14), (19) and the assumption we have
which reveals the conclusion of Lemma 7.
β
Lemma 8.
Suppose that and Then
Proof.
Since are non-constant meromorphic functions satisfying (4), we deduce by (11)- (13) and the assumption of Lemma 8 that there exists only one of that is equal to zero, the other two of are negative integers, say and This together with (10) and Lemma 1(v) gives
By (14), (21) and the assumption we have
which reveals the conclusion of Lemma 8.
β
Lemma 9.
Suppose that and
Then
Proof.
Since are non-constant meromorphic functions satisfying (4), we deduce by (11)- (13) and
the assumption
that This together with (10) and Lemma 1(iii) gives
By (14) and (23) we have
which reveals the conclusion of Lemma 9.
β
Lemma 10.
Suppose that and
If then
Proof.
Since are non-constant meromorphic functions satisfying (4), we deduce by (11)-(13) and
the assumption and
that This together with (10) and Lemma 1(iv) gives
By (14) and (25) we have
which reveals the conclusion of Lemma 10.
β
Let be three non-constant meromorphic functions. Next we denote by
the reduced counting function of common poles of and in and denote by the reduced
counting function of common zeros of and in
The following result plays an important role in proving Theorem 8 of this paper:
Lemma 11.
Suppose that there exist non-constant meromorphic functions
and satisfying the functional equation (1) for Then
and
where and are defined as in Lemma 2.
Proof.
We first set
Since are non-constant meromorphic functions satisfying (1) for by Lemma 5 we have
where is the Wronskian determinant of and and is defined as in Lemma 2. Moreover, (14) can be rewritten as
By (29) and Lemmas 6-10 we deduce
From (28)-(30) we have
which implies that
and
By (31)-(37) we get the conclusion of Lemma 11.
β
In the same manner as in the proof of Lemma 11 we can get the following two results that is used to prove Theorem 9:
Lemma 12.
Suppose that there exist non-constant entire functions
and satisfying the functional equation (1) for Then
and
where and are defined as in Lemma 2.
Next we introduce some notions of normal families of meromorphic functions in a domain of the complex plane. Following Montel [27], a class of meromorphic functions in a domain is called normal in if given any sequence of meromorphic functions in we can find a subsequence which converges everywhere in and uniformly on compact subsets of with respect to the chordal metric on Riemann sphere. We then say that converges locally uniformly in
An equivalent statement is that for every in there exists a neighbourhood in which or converges uniformly as Suppose in fact that the second condition is satisfied. If are two points in the -plane, their distance in the chordal metric of Riemann sphere is
Also
Thus, if either or uniformly in a set then
uniformly in Thus, if one of these two conditions holds uniformly in some neighbourhood of every point of
then uniformly in some neighbourhood of every point of and hence by the Heine-Borel theorem uniformly on every compact subset of see, for example, [18, pp.157-160]. We recall the following classical normality criterion of a class of meromorphic functions due to Marty[25]:
Lemma 13.
([25]) A class of functions meromorphic in a domain of the complex plane is normal in if and only if is uniformly bounded on any compact subset of for
The following result is the well-known Pang-Zalcmanβs Lemma:
Lemma 14.
(Pang-Zalcmanβs Lemma, [31, 43]).
Let be a family of meromorphic functions in the unit disc and be a real number
satisfying Then, if is not normal at a point
there exist, for each :
(i) β points
(ii) β positive numbers
and
(iii) β functions such that spherically uniformly on compact subset of
where is a non-constant meromorphic function. The function
may be taken to satisfy the normalization
Next we follow Bergweiler [1, p.153] to introduce some definitions and results in the theory of iterations of the meromorphic functions in the complex plane that play an important role in proving the main results in this paper: let be a meromorphic
function, where is the complex plane, and we assume that is neither a constant nor a linear
transformation. We denote by the -th iterate of that is, with and
with for . Then is defined for all except
for a countable set which consists of the poles of If is
rational, then has a meromorphic extension to and we denote the extension
again by we see that is defined and meromorphic in But if
is transcendental in the complex plane, it is not reasonable for us to define
The basic objects studied in iteration theory are the Fatou set and
the Julia set of a meromorphic function Roughly speaking, the
Fatou set is the set where the iterative behavior is relatively tame in the sense
that points close to each other behave similarly, while the Julia set is the set where chaotic phenomena take place. The formal definitions are
and where and in what follows, We also need the following notion of the periodic point in iteration theory for proving the main results in this paper:
Definition 1.
By definition, is called a periodic point of a non-constant meromorphic function if for
some positive integer satisfying In this case, the positive integer is called a period of the point and the smallest positive integer with this property is called the minimal period of the point For a periodic point 0f
minimal period is called the multiplier of the point We mention that which
can happen only for the rational function of course, this has to be modified. In this case, the multiplier is defined to be
where A periodic point is called attracting, indifferent, or repelling accordingly as the
modulus of its multiplier is less than, equal to, or greater than Periodic
points of multiplier are called super attracting. Some writers reserve the
term attracting for the case but we consider super attracting
as a special case of attracting. The multiplier of an indifferent periodic point is of the form where We say that is rationally indifferent if is rational and irrationally indifferent otherwise. Also, a point is called
pre-periodic if is periodic for some Finally, a periodic point of
period is called a fixed point. It is easy to see that attracting periodic points are in while repelling
and rationally indifferent periodic points are in For irrationally indifferent
periodic points the question whether they are in or is difficult to decide.
Both possibilities do occur. We refer the reader to the classical papers of Cremer
[4, 5] and Siegel [33], as well as more recent work of Yoccoz [41, 42]. An
exposition of these and other results, together with further references, can be
found in [24]. The behavior of the iterates in the neighborhood of a fixed point (or, more
generally, a periodic point) is intimately connected with the solution of certain
functional equations. For most results in this direction, it is required only that
the function under consideration is defined in a neighborhood of the fixed point,
and it is usually irrelevant whether it extends to a rational or transcendental
meromorphic function. Therefore, we omit this topic here but refer to the
pβapers and books on iteration of rational functions cited in the introduction.
We recall the following results from Bergweiler [1, p.153]:
Lemma 15.
([1, p.161, Theorem 5]) If is a transcendental meromorphic function in the complex plane and is a positive integer such that then has infinitely many repelling periodic points of minimal period
Lemma 16.
([1, p.160, Theorem 4]) Let f be a non-constant meromorphic function in the complex plane. Then, the Julia set of the meromorphic function is the closure of the set of repelling periodic points of f.
The following result is due to Clunie [3]:
Lemma 17.
([3]) Let be a transcendental meromorphic function in the complex plane, and let be a non-constant entire function. Then
3. Notations and definitions for the proofs
Before we prove the main results of this paper, we first introduce the following notations and definitions: Let and be two non-constant meromorphic functions in the complex plane, and let be a value in the extended plane. Let
βcountβ those points in where is taken by and with the same
multiplicity, and each point is counted only once, and let
be the reduced counting function of the
common -points of and in
where means We
say that and share the value CMβ, if
These definitions can be found in [35], where CMβ here is denoted by βCMβ in [35]. We also need the following definition:
Definition 2.
([23, Definition 1]) Let be a positive integer and Next we denote by the counting function of those -points of (counted with proper multiplicities) whose multiplicities are not greater than and denote by the counting function of those -points of (counted with proper multiplicities) whose multiplicities are not less than We denote by and the reduced forms of and respectively. Here and
mean and respectively.
Definition 3.
([13, Definition 2.1])
For a meromorphic function satisfying and a positive
integer let denote the number of zeros of in counted in
the following manner: a zero of of multiplicity is counted exactly times where
Then let denote the corresponding integrated counting
function; that is,
Regarding the well-known integrated counting functions and
we see that and
Secondly, we need generalizations of previous definitions. Let be non-constant meromorphic functions. We denote by the reduced counting function of common poles of in and denote by
the reduced counting function of common zeros of in Restrictions can be given in these definitions. For instance, the notation will mean the reduced counting function of common zeros of and in that are not zeros of The notation will mean the reduced counting function of common zeros of and in that are not zeros of and In addition, the notation
denotes the counting function of those zeros of in that are zeros of where each point in
is counted according to its
multiplicity as a zero of The notation
denotes the counting function of those zeros of in that are poles of where each point in
is counted according to its multiplicity as a zero of We denote by and the reduced forms of and respectively.
The notation denotes the reduced counting function of those simple zeros of in that are zeros of
The notation denotes the reduced counting function of those simple zeros of in that are poles of
Thirdly, we denote by the reduced counting function of common simple poles of and in denote by the reduced counting function of common simple zeros of and in denote by the reduced counting function of those zeros of in that are not zeros of and denote by the reduced counting function of those zeros of in that are neither zeros of nor zeros of
Last, we denote by the reduced counting function of those common simple zeros in where each such common simple zero of and is neither a zero of nor a zero of
4. Proof of Theorem 8
Suppose that there exist three transcendental meromorphic functions and
satisfying (1) for
We consider the following two cases:
Case 1. β Suppose that and are linearly dependent. Then there exist three constants satisfying such that
Without loss of generality, we suppose that Then, by substituting (40) into we have
By (41) and the well-known result from Gross [9] that the equation cannot have non-constant meromorphic solutions and we get a contradiction.
Case 2. β Suppose that and are linearly independent. First of all, by Lemma 11 we can deduce that and share and CM* such that
and
where and in what follows, and are defined as in Lemma 2.
Also by Lemma 11 we have
which means that any two of the functions and almost have no common zeros. Moreover, by (42)-(44) we deduce
and
On the other hand, by (4), (10) and Theorem 5 we have for that
where is a small function of and such that
On the other hand, by taking the first derivatives of two sides of
(1) for we have
and so we have
and
Now we prove the following claim:
Claim I. The following items hold: (i) and share CM*; (ii) and share CM*; (iii) and share CM*.
Without loss of generality, next we only prove Claim I (i): First of all, by (42) we have
By (51) we have
Next we prove
For this purpose, now we let be a common zero of and in the complex plane. Then at least one of the following four cases must hold: or We discuss these four cases as follows:
(1). Suppose that Then, by (47) we can deduce that the multiplicity of as a zero of
is not less than We discuss this as follows:
Suppose that and Then, by comparing the multiplicities of as the zeros of two sides of (54) we deduce that is a zero of
Therefore,
and so
Suppose that or Then, by the supposition we can see that is a common zero of and or a common zero of and Moreover, by (43) and (44) we have
and
and so
By (56) and (59) we have
(2). Suppose that Then, by (48) we have and so either or We discuss this as follows:
Suppose that Then, by (47) and the supposition we deduce that the multiplicity of as a zero of
and
is not less than This implies that, possibly is a multiple zero of or a zero of one of and or a zero of Therefore, by (43) and (44), we deduce
and so
Suppose that Then, we deduce by (47) and the supposition that possibly is a multiple zero of one of and or a zero of Therefore, by (43) we deduce
and so
By (63), (64) and the above analysis we have
(3). Suppose that Then from (44) we have
(4). Suppose that holds. Then, in the same manner as in the above case (1) we deduce by (43), (44) and (47) that
and
By (67) and (68) we have
Finally, by (60), (65), (66), (69) and the above analysis in cases (1)-(4) we have
which is (53). By (49), (52) and (53) we have Claim I(i). Next we prove the following claim:
Claim II. β The following equalities hold:
where and in what follows, denotes the counting function of those zeros in that are not zeros of and have similar definitions.
Indeed, by Claim I(i) we know that and share CM*, and by (43) we know that almost all the zeros of and in the complex plane are simple zeros such that
By (71) we have
and
Let be a common zero of and that is neither a zero of nor a zero of and that is a simple zero of at least. By (47) we have (61) and (62). This is discussed as follows:
Suppose that Then, by (61) we deduce
Suppose that is a simple zero of Then, by the supposition we deduce that the multiplicity of as a zero of is equal to at least. By (61) and (62) we have
Therefore, the multiplicity of as a zero of is also equal to at least. This together with (61) and the supposition that is a simple zero of implies that is a zero of Therefore, by (61) we deduce
By (75) and (76) we have
By Claim I(i) we know that and share CM*. Combining this with (42)-(44), (61), (71)-(77) and the second fundamental theorem, we have
By (78) we obtain that all the inequalities in (78) are equalities that are equal to This gives
and
By Claim I(ii) we know that and share CM*. Next, in the same manner as the above, we can deduce by (42)-(44), (54), (71)-(74) and the second fundamental theorem that
By (82) we obtain that all the inequalities in (82) are equalities that are equal to This gives
and
Obviously
and
By (79)-(81) and (83)-(87) we deduce Claim II(v)-Claim II(vii). This together with (42), (43), (78) and (82) gives Claim II(i)-Claim II(iv). This proves Claim II.
Noting that and are transcendental meromorphic functions in the complex plane, we use Lemma 15 to deduce that has infinitely many repelling periodic points of minimal period for a given positive integer such that Next we let be a repelling periodic point of minimal period Then, it follows from Lemma 16 that with
and
for each positive integer Here and in what follows,
and is defined as
for each positive integer Here is some small positive number such that is analytic in for each positive integer and such that
and
for each and each positive integer This is derived from (88)-(90) and the continuity of and at From (91) we have
for each positive integer Next we let
for each positive integer Then, it follows from (94) and (95) that
for each positive integer From (88)-(92) and (96) we have
for each positive integer and such that for each positive integer we have
with and Noting that we deduce by (90), (91) and the definition of the Julia set of the transcendental meromorphic function that the family with for each positive integer is not normal at the point Accordingly, it follows from (96) and (97) that the family defined in is not normal at the origin point Combining this with Lemma 14, we see that there exist (i) points in such that (ii) positive numbers such that
and (iii) an infinite subsequence of say itself
such that
spherically uniformly on compact subset of Here and in what follows,
for each and the large positive integer and is a non-constant meromorphic function. The function
may be taken to satisfy the normalization Moreover, from (93) and (100) we have
for each and the large positive integer such that with and Next we prove the following claim:
Claim III. Under the assumptions of Theorem 8, the function defined in (99) is a transcendental entire function such that for each Here and in what follows, denotes the first derivative of the composition of and
We prove Claim III: β first of all, we suppose that there exists a point such that Then, there exists some positive number such that is analytic in and such that
for each Combining this with the known result
spherically uniformly on the closed disk where
and with are analytic functions on and is the only one zero of in by Hurwitzβs theorem (cf.[32, p.9 Hurwitz Theorem]) we deduce that has one zero in at least for the large positive integer However, from (99), (100), the left inequality of (98) and the supposition we deduce that
has no zeros in This is a contradiction. Therefore, we prove that
is a non-constant entire function.
Secondly, we suppose that there exists a point such that Then, there exists some positive number such that is analytic in and such that
for each
Combining this with the known result that spherically uniformly on the closed disk where and with are analytic functions on and is the only one zero of in we deduce by Hurwitzβs theorem (cf.[32, p.9 Hurwitz Theorem]) that has one zero in at least for the large positive integer However, from (99), (100), the left inequality of (98) and the supposition we deduce that
has no zeros in for the large positive integer m. This is a contradiction. This implies that is a Picard exceptional value of in the complex plane. Combining this with the fundamental theorem of algebra and the known result that is a non-constant entire function, we deduce that is a transcendental entire function such that is a Picard exceptional value of
Thirdly, we prove that is a Picard exceptional value of Indeed, from the known result that is a transcendental entire function and the supposition that is a transcendental meromorphic function in the complex plane, we deduce from Lemma 17 that the composition of and is also a transcendental meromorphic function in the complex plane. Then,
the first derivative of the composition is still a transcendental meromorphic function in the complex plane. Next we suppose that there exists a point such that Then, we can find some positive number such that is analytic in and such that
for each On the other hand, from (99) and (100), we deduce that
spherically uniformly on the closed disk
Noting that is analytic on the closed disk we see that the
the composition is also analytic on the closed disk Combining this with the supposition that the composition is analytic on the closed disk
for the large positive integer we deduce by (102) and Weierstrass theorem (cf.[32, p.9, Weierstrass Theorem]) that
spherically uniformly on the closed disk On the other hand, since
is analytic on the closed disk such that for each and is the only one zero of on the closed disk we deduce by (103), Hurwitzβs theorem (cf.[32, p.9 Hurwitz Theorem]) and the supposition that the composition is analytic on the closed disk
for each positive integer we deduce that
has one zero at least in the open disk
However, from (93)-(96), (99)-(101) and the right formula of (98) we deduce
for any given and the corresponding large positive integer such that (100) holds. From (105) we see that (104) has no zeros on the closed disk for the corresponding large positive integer This is a contradiction. Therefore, we prove that is a Picard exceptional value of This complete the proof of Claim III.
Last, we complete the proof of Theorem 8. Indeed, from Claim III we know that in (99) is a transcendental entire function. Combining this with the supposition that the transcendental meromorphic functions satisfy the functional equation we deduce that
and are still transcendental meromorphic functions that satisfy
for each From (106) and Claim II(V)-Claim II(vii) we have
where and in what follows, is any quantity such that as and Here is a set of finite linear measure, and From (106) and Lemma 11 we have
From Claim III we have
We substituting (109) into (107), and then we have
which contradicts (108). Theorem 8 is thus completely proved.
5. Proof of Theorem 9
Since the proof is similar to the proof of Theorem 8, we omit some details below.
Case 1. Suppose that and are linearly dependent. Then, in the same manner as in Case 1 in the proof of Theorem 8 we can get a contradiction.
Case 2. Suppose that and are linearly independent. First of all, by Lemma 12 we can deduce that and share CM* such that (43) holds. Also by Lemma 12 we have (44). By (44) we can see that
any two of the functions and almost have no common zeros. Moreover, by (43) we deduce
and
On the other hand, by (4), (10) and Theorem 6 we have for that
where is a small function of and such that
On the other hand, by taking the first derivatives of two sides of
(1) for we have
and so we have
and
In the same manner as in the proof of Claim I and by (43), (44) and (110)-(115) we get the following claim:
Claim IV. The following items hold: (i) and share CM*; (ii) and share CM*; (iii) and share CM*.
Next we prove the following claim:
Claim V. β The following equalities hold:
Indeed, by Claim IV(i) we know that and share CM*, and by (110) we know that almost all the zeros of and in the complex plane are simple zeros such that
By (116) we have
and
On the other hand, by (112) we have
and
Now we let be a common zero of and that is neither a zero of nor a zero of and that is a simple zero of at most. This is discussed as follows:
Suppose that Then, by (120) and the supposition we deduce that is a zero of such that
Suppose that is a simple zero of Then, by the supposition we deduce that the multiplicity of as a zero of is equal to at least. By (120) and (122) we have
Therefore, the multiplicity of as a zero of is also equal to at least. This together with (120) and the supposition that is a simple zero of implies that is a zero of Therefore, by (120) we deduce
By (123) and (124) we have
By Claim IV(i) we know that and share CM*. Combining this with (43), (44), (112), (116)-(119), (125) and the second fundamental theorem, we have
By (126) we obtain that all the inequalities in (126) are equalities that are equal to This gives
and
By Claim IV(ii) we know that and share CM*. Next, in the same manner as the above, we can deduce by (43), (44), (112), (116)-(125) and the second fundamental theorem that
By (130) we obtain that all the inequalities in (130) are equalities that are equal to This gives
and
Obviously
and
By (127)-(129) and (131)-(135) we deduce Claim V(v)- Claim V(vii). This together with (43), (126) and (130) gives Claim V(i)- Claim V(iv).
From Lemma 15 and the supposition that and are transcendental entire functions, we deduce that has infinitely many repelling periodic points of minimal period for the given positive integer such that As in the proof of Theorem 8, next we let be a repelling periodic point of minimal period and we have from Lemma 16 that As in the proof of Theorem 8, we have (88)-(98). Noting that we deduce by (90), (91) and the definition of the Julia set of the transcendental entire function that the family with for each positive integer is not normal at the point Accordingly, it follows from (96) and (97) that the family defined in is not normal at the origin point Combining this with Lemma 14, we see that there exist (iv) points (v) positive numbers
and (vi) an infinite subsequence of say itself
such that (99) holds, and such that (100) holds for each and the large positive integer Here is a non-constant entire function. The function
may be taken to satisfy the normalization Moreover, from (93) and (100) we have (101) for each and the large positive integer such that with and Next we use (98)-(101) and the lines of the reasoning of the proof of Claim III in the proof of Theorem 8 to get the following claim:
Claim VI. β Under the assumptions of Theorem 9, the function defined in (99) is a transcendental entire function such that for each Here and in what follows, denotes the first derivative of the composition of the entire functions and
Last, we complete the proof of Theorem 9. Indeed, from Claim VI we know that in (99) is a transcendental entire function. Combining this with the supposition that the transcendental entire functions satisfy the functional equation we deduce that
and are still transcendental entire functions that satisfy
for each From (136) and Claim V(V)-Claim V(vii) we have
where and in what follows, is any quantity such that as and Here is a set of finite linear measure, and From (136) and Lemma 12 we have
From Claim VI we have
We substituting (139) into (137), and then we have
which contradicts (138). Theorem 9 is thus completely proved.

### v5

@@
 non-constant meromorphic function we denote by any quantity
-satisfying as where is the Nevanlinna characteristic of Following [17, p.466], we define the notion of the small function of a non-constant meromorphic function in the complex plane as follows: let be a non-constant meromorphic function.
+satisfying as where is the Nevanlinna characteristic of Following [18, p.466], we define the notion of the small function of a non-constant meromorphic function in the complex plane as follows: let be a non-constant meromorphic function.
 and let be also a meromorphic function in the complex plane such that for all possibly outside a set of finite linear measure.
@@
 Theorem 5.
-([21, Proposition 5.1]).
+([22, Proposition 5.1]).
 Suppose that there exist transcendental meromorphic functions
@@
 Theorem 6.
-([21, Proposition 6.1]).
+([22, Proposition 6.1]).
 Suppose that there exist transcendental functions
@@
 with respect to and such that
-For the existence of non-constant meromorphic solutions of a much more general three term functional equation, we recall the following result from [20] and Yang-Zhang [37, 38]:
+For the existence of non-constant meromorphic solutions of a much more general three term functional equation, we recall the following result from Hu-Li-Yang [21] and Yang-Zhang [39, 40]:
 Theorem 7.
@@
 uniformly in Thus, if one of these two conditions holds uniformly in some neighbourhood of every point of
-then uniformly in some neighbourhood of every point of and hence by the Heine-Borel theorem uniformly on every compact subset of see, for example, [18, pp.157-160].
-We recall the following classical normality criterion of a class of meromorphic functions due to Marty[25]:
+then uniformly in some neighbourhood of every point of and hence by the Heine-Borel theorem uniformly on every compact subset of see, for example, [19, pp.157-160].
+In addition, we need the following notion of the spherical derivative of a meromorphic function in the complex plane (cf.[19, 38]: let be a non-constant meromorphic function.
+The spherical derivative of
+at is given as We recall the following classical normality criterion of a class of meromorphic functions due to Marty[26]:
 Lemma 13.
 ([26]) A class of functions meromorphic in a domain of the complex plane is normal in if and only if is uniformly bounded on any compact subset of for
+The following result is due to Pang [33]:
+Lemma 14.
+([33, Lemma 1])
+Let be a meromorphic function in the domain and let be a real number satisfying If there exists a point where is a fixed positive number satisfying such that
+then there exist a point and a real number satisfying such that
 The following result is the well-known Pang-Zalcmanβs Lemma:
 Lemma 15.
-(Pang-Zalcmanβs Lemma, [31, 43]).
+(Pang-Zalcmanβs Lemma, [17, 32, 45]).
 Let be a family of meromorphic functions in the unit disc and be a real number
 satisfying Then, if is not normal at a point
+there exist, for each
+(i) points
+(ii) positive numbers
+and
+(iii) functions such that spherically uniformly on compact subset of
+where is a non-constant meromorphic function. The function
+may be taken to satisfy the normalization
+From Remark 2 (III) and the proof of Lemma 15 for in Remark 2 (II), we get the following result:
+Lemma 16.
+Let be a family of holomorphic functions in a disc where is a point, is a domain in the complex plane, and is an infinite sequence of positive numbers such that for each positive integer and let be a real number
+satisfying Then, if is not normal at the point
 there exist, for each :
@@
 and
-(iii) β functions such that spherically uniformly on compact subset of
-where is a non-constant meromorphic function. The function
+(iii) β an infinite subsequence of the family say itself, such that spherically uniformly on compact subset of
+where is a non-constant entire function. The function
 may be taken to satisfy the normalization
@@
 Definition 2.
-([23, Definition 1]) Let be a positive integer and Next we denote by the counting function of those -points of (counted with proper multiplicities) whose multiplicities are not greater than and denote by the counting function of those -points of (counted with proper multiplicities) whose multiplicities are not less than We denote by and the reduced forms of and respectively.
+([24, Definition 1]) Let be a positive integer and Next we denote by the counting function of those -points of (counted with proper multiplicities) whose multiplicities are not greater than and denote by the counting function of those -points of (counted with proper multiplicities) whose multiplicities are not less than We denote by and the reduced forms of and respectively.
 Here and
@@
 the reduced counting function of common zeros of in Restrictions can be given in these definitions.
-For instance, the notation will mean the reduced counting function of common zeros of and in that are not zeros of The notation will mean the reduced counting function of common zeros of and in that are not zeros of and In addition, the notation
-denotes the counting function of those zeros of in that are zeros of where each point in
-is counted according to its
-multiplicity as a zero of The notation
-denotes the counting function of those zeros of in that are poles of where each point in
-is counted according to its multiplicity as a zero of We denote by and the reduced forms of and respectively.
-The notation denotes the reduced counting function of those simple zeros of in that are zeros of
-The notation denotes the reduced counting function of those simple zeros of in that are poles of
-Thirdly, we denote by the reduced counting function of common simple poles of and in denote by the reduced counting function of common simple zeros of and in denote by the reduced counting function of those zeros of in that are not zeros of and denote by the reduced counting function of those zeros of in that are neither zeros of nor zeros of
-Last, we denote by the reduced counting function of those common simple zeros in where each such common simple zero of and is neither a zero of nor a zero of
+For instance, the notation will mean the reduced counting function of common zeros of and in that are not zeros of The notation will mean the reduced counting function of common zeros of and in that are not zeros of and The notation will mean the reduced counting function of common zeros of in that are either zeros of or zeros of
+Thirdly, we denote by the reduced counting function of common simple poles of and in and denote by the reduced counting function of common simple zeros of and in
 4. Proof of Theorem 8
@@
 and is defined as
-for each positive integer Here is some small positive number such that is analytic in for each positive integer and such that
+for each positive integer Here is some small positive number satisfying for each positive integer such that is analytic in for each positive integer and such that
 and
@@
 spherically uniformly on the closed disk where
-and with are analytic functions on and is the only one zero of in by Hurwitzβs theorem (cf.[32, p.9 Hurwitz Theorem]) we deduce that has one zero in at least for the large positive integer However, from (99), (100), the left inequality of (98) and the supposition we deduce that
+and with are analytic functions on and is the only one zero of in by Hurwitzβs theorem (cf.[34, p.9 Hurwitz Theorem]) we deduce that has one zero in at least for the large positive integer However, from (123), (124), the left inequality of (122) and the supposition we deduce that
 has no zeros in This is a contradiction. Therefore, we prove that
@@
 for each
-Combining this with the known result that spherically uniformly on the closed disk where and with are analytic functions on and is the only one zero of in we deduce by Hurwitzβs theorem (cf.[32, p.9 Hurwitz Theorem]) that has one zero in at least for the large positive integer However, from (99), (100), the left inequality of (98) and the supposition we deduce that
+Combining this with the known result that spherically uniformly on the closed disk where and with are analytic functions on and is the only one zero of in we deduce by Hurwitzβs theorem (cf.[34, p.9 Hurwitz Theorem]) that has one zero in at least for the large positive integer However, from (123), (124), the left inequality of (122) and the supposition we deduce that
 has no zeros in for the large positive integer m.
@@
 the composition is also analytic on the closed disk Combining this with the supposition that the composition is analytic on the closed disk
-for the large positive integer we deduce by (102) and Weierstrass theorem (cf.[32, p.9, Weierstrass Theorem]) that
+for the large positive integer we deduce by (126) and Weierstrass theorem (cf.[34, p.9, Weierstrass Theorem]) that
 spherically uniformly on the closed disk On the other hand, since
-is analytic on the closed disk such that for each and is the only one zero of on the closed disk we deduce by (103), Hurwitzβs theorem (cf.[32, p.9 Hurwitz Theorem]) and the supposition that the composition is analytic on the closed disk
+is analytic on the closed disk such that for each and is the only one zero of on the closed disk we deduce by (127), Hurwitzβs theorem (cf.[34, p.9 Hurwitz Theorem]) and the supposition that the composition is analytic on the closed disk
 for each positive integer we deduce that
@@
 By (151)-(153) and (155)-(159) we deduce Claim V(v)- Claim V(vii). This together with (67), (150) and (154) gives Claim V(i)- Claim V(iv).
-From Lemma 15 and the supposition that and are transcendental entire functions, we deduce that has infinitely many repelling periodic points of minimal period for the given positive integer such that As in the proof of Theorem 8, next we let be a repelling periodic point of minimal period and we have from Lemma 16 that As in the proof of Theorem 8, we have (88)-(98).
-Noting that we deduce by (90), (91) and the definition of the Julia set of the transcendental entire function that the family with for each positive integer is not normal at the point Accordingly, it follows from (96) and (97) that the family defined in is not normal at the origin point Combining this with Lemma 14, we see that there exist (iv) points (v) positive numbers
+From Lemma 17 and the supposition that and are transcendental entire functions, we deduce that has infinitely many repelling periodic points of minimal period for the given positive integer such that As in the proof of Theorem 8, next we let be a repelling periodic point of minimal period and we have from Lemma 18 that As in the proof of Theorem 8, we have (112)-(122), where and in what follows, is some small positive number satisfying for each positive integer Noting that we deduce by (114), (115) and the definition of the Julia set of the transcendental entire function that the family with for each positive integer is not normal at the point Accordingly, it follows from (120) and (121) that the family defined in is not normal at the origin point Combining this with Lemma 16, we see that there exist (iv) points in such that (v) positive numbers such that
 and (vi) an infinite subsequence of say itself